OpenStudy (gabylovesyou):

Choose the equation below that represents the line passing through the point (1, -4) with a slope of 1/2.

OpenStudy (gabylovesyou):

y - 4 = 1/2(x - 1) y + 4 = 2(x - 1) y + 4 = 1/2(x - 1) y - 4 = 2(x + 1)

OpenStudy (gabylovesyou):

@telltoamit

OpenStudy (anonymous):

hmm

OpenStudy (anonymous):

the equation that represents the line passing through the point (x1, y1) with a slope of m: y - y1 = m (x - x1)

OpenStudy (gabylovesyou):

c?

OpenStudy (anonymous):

yepp, brilliant!

OpenStudy (anonymous):

good job